OpenStudy (anonymous):

what the domain and range for this y=log5x

OpenStudy (turingtest):

for any log function\[y=\log_ax\iff a^y=x\]notice that the second statement means that x can never be zero or negative

OpenStudy (anonymous):

so what the domain and range

OpenStudy (turingtest):

for example\[\log_24=2\iff2^2=4\]\[\log_2(\frac12)=-1\iff2^{-1}=\frac12\]\[\log_2(\frac14)=-2\iff2^{-2}=\frac14\]now what about\[\log_20\]? No matter what power we raise 2 to, we will never get zero. We can only get very small positive numbers by raising 2 to very large negative numbers. Hence \(\log x\) is undefined for \(x\le0\) so the domain is \(\left\{ x|x>0 \right\}\)

OpenStudy (anonymous):

can u tell me the domain and range please

OpenStudy (turingtest):

Please actually read what I wrote I gave the domain above Try to figure out the range based on what I told you about the domain Welcome to Open Study :)